OpenStudy (anonymous):

a jar contains 24 marbels,some are green others are blue . if a marble is drawn at random from the jar the probability it is green is 2/3 .find the number of blue and green marbles in the jar

OpenStudy (anonymous):

number of marbles = 24 number of green balls = x = x/24 2/3 = x/24 2*24/3 = x x = 16 number of blue and green marbles = 24 number of blue marbles + 16 = 24 How many blue and green marbles are there?

OpenStudy (anonymous):

two thirds of the marbles are green, one third of them are blue

OpenStudy (anonymous):

@donmath

OpenStudy (anonymous):

hahha

OpenStudy (anonymous):



OpenStudy (anonymous):

It means u should take the total jars=M

OpenStudy (anonymous):